OpenStudy (anonymous):

if a car is moving at 13m/s and accelerates at constant rate of 3.0m/s^2 for 5 sec, how fast is it then going?

OpenStudy (anonymous):

use V=at+Vi

OpenStudy (jhannybean):

Use kinematics to solve this. Given: initial velocity = 13 m/s acceleration = 3.0 m/s^2 time = 5 sec use \[\large V _{f}=V _{i}+ a _{c}t\]\

OpenStudy (shamim):

plz c my attachment